# Reinforcement learning with combinatorial actions for coupled restless bandits

### v1

@@ Abstract
Reinforcement learning (RL) has increasingly been applied to solve real-world planning problems, with progress in handling large state spaces and time horizons. However, a key bottleneck in many domains is that RL methods cannot accommodate large, combinatorially structured action spaces. In such settings, even representing the set of feasible actions at a single step may require a complex discrete optimization formulation. We leverage recent advances in embedding trained neural networks into optimization problems to propose SEQUOIA, an RL algorithm that directly optimizes for long-term reward over the feasible action space. Our approach embeds a Q-network into a mixed-integer program to select a combinatorial action in each timestep. Here, we focus on planning over restless bandits, a class of planning problems which capture many real-world examples of sequential decision making. We introduce coRMAB, a broader class of restless bandits with combinatorial actions that cannot be decoupled across the arms of the restless bandit, requiring direct solving over the joint, exponentially large action space. We empirically validate SEQUOIA on four novel restless bandit problems with combinatorial constraints: multiple interventions, path constraints, bipartite matching, and capacity constraints. Our approach significantly outperforms existing methods—which cannot address sequential planning and combinatorial selection simultaneously—by an average of 26.4% on these difficult instances.
@@ 1 Introduction
Reinforcement learning (RL) has made tremendous progress in recent years to solve a wide range of practical problems Treloar et al. (2020); Marot et al. (2021); Silvestro et al. (2022); Degrave et al. (2022). While successful at dealing with large or infinite state spaces, RL is limited when the action space is both combinatorial and discrete. This limitation is pertinent to many real-world sequential decision-making problems, where resource constraints frequently lead to combinatorial action spaces Dulac-Arnold et al. (2020). Consider, for example, a sequential resource allocation problem in which public health workers are dispatched to visit patients. The workers each have a limited daily budget to maximize patient well-being. These requirements give rise to an exponentially large combinatorial action space to optimize over, even when the number of workers and patients is small.
@@
A suitable framework for modeling such problems is that of restless multi-armed bandits Niño-Mora (2023), which have been applied to settings from clinical trial design Villar et al. (2015) to patient interventions Mate et al. (2022). An “arm” of the restless bandit corresponds to a patient in the aforementioned public health example. At every timestep, each arm is in one of a finite number of states (e.g., representing the health of that individual), and we aim to select actions so that arms move into the more beneficial states. Typically, restless bandit solutions have been confined to relatively simple and thus tractable problem structures. For example, the common Whittle index policy Whittle (1988) requires that each arm be independent, each action impacts only a single arm, and arms transition independently (Figure 1A). Under these assumptions, the problem admits a threshold-based top-$K$ policy which decouples the combinatorial action selection by learning policies for each arm independently.
@@
Figure 1: Restless bandits are a variant of planning over Markov decision processes, where each “arm” transitions depending on whether it is acted upon. Standard restless bandits can be solved by threshold-based policies, but these approaches are unable to address more challenging settings with combinatorial constraints on the actions that cannot be decoupled, prohibiting the application of easy heuristic solutions. Our paper considers this class of strongly coupled restless bandits (coRMAB). We describe these new problem formulations (AD) for restless bandits in detail in Section 3.
@@
Real-world applications, on the other hand, often involve complicated problem structures, beyond the simple budget constraint, for which action decomposition would substantially degrade performance; Figure 1A–D illustrate four such scenarios. For example, consider the problem of planning a route for a health worker to visit patients within a travel constraint (see Figure 1D): each action (path) impacts multiple patients and cannot be decomposed further (e.g., into separate edges) while ensuring the overall path is feasible. Another example is where actions represent heterogeneous interventions that may impact not just a single arm but several simultaneously. Individuals may be acted on by multiple actions simultaneously (e.g., exposure to different messaging campaigns or nurse visits), with a potentially nonlinear cumulative effect (Figure 1B).
@@
Here, we consider for the first time sequential combinatorial settings where the reward comes not from a single-shot action but is incurred after enacting a policy over many timesteps. We leverage recent advances in integrating deep learning with mathematical programming, in which a neural network with ReLU activations can be directly embedded into a mixed-integer program. Incorporated in an RL training procedure, this technique enables our approach to learn the long-term reward using a deep learning approximation of the value function, while tractably optimizing this function over the discrete combinatorial space at each step with a mixed-integer program. We call our algorithm SEQUOIA, for SEQUential cOmbInatorial Actions.
@@ 2 Problem description
We consider offline, stochastic planning for restless bandits with combinatorial action constraints, where the transition dynamics and reward are known a priori, but the state and action space are too large to be solved directly. Specifically, we consider the setting where the arms are strongly coupled so we cannot decouple the arms of the bandit into independently controlled Markov chains using a Lagrangian relaxation. We begin by describing the standard restless bandit, then present a general formulation for constrained combinatorial-action RMABs, which we refer to as coRMAB.
@@ 2.2 Restless bandits with strongly coupled actions (coRMAB)
We now introduce coRMABs, a broader class of problems for restless bandits where actions may be strongly coupled due to combinatorial constraints. We consider a set of $N$ actions $a_{i}$ that each impact the transition probability of some subset of the arms. At each timestep, we must pick a combinatorial action over the arms, i.e., the action vector $\bm{a}$ must be in $\mathcal{C}\subseteq\mathcal{A}^{\times}$, a set defined by constraints, instantiations of which we provide in Section 3. Now the transition probabilities may be coupled, with individual actions impacting multiple arms. We denote the joint transition probability as $P^{\times}:\mathcal{S}^{J}\times\mathcal{A}^{N}\times\mathcal{S}^{J}\rightarrow[0,1]^{J}$.
@@ 4 Solving RMABs with combinatorial actions
We present a novel algorithm, SEQUOIA, which builds on deep Q-learning (DQN) by integrating a mixed-integer linear program (MILP) for combinatorial action selection. We start by introducing the basic DQN+MILP framework in Section 4.1. Then, in Section 4.2, we provide a set of strategies to find a good initialization for the Q-network and improve computational efficiency. We visualize our approach in Figure 2, and Algorithm 1 provides pseudocode for the Q-network training procedure.
While we focus on the restless bandit problem setting, SEQUOIA generalizes to other sequential planning problems with combinatorial actions, provided that the restrictions on the actions can be represented as constraints in a mixed-integer program.
@@ 4.1 Q-learning with combinatorial actions
Deep Q-learning aims to estimate the long-run value of action $\bm{a}$ in state $\bm{s}$, denoted $Q(\bm{s},\bm{a})$, by parameterizing the Q-function with a neural network. The DQN algorithm Mnih (2013) samples (action, next state, reward) sequences via a mix of random actions and on-policy samples (i.e., actions that maximize the current estimate of $Q$). The samples are used to improve the Q-function via temporal difference (TD) updates which aim to minimize the residual in the Bellman equation.
@@
Instead, we leverage recent advances that enable encoding a neural network into a mixed-integer linear program to enable optimization Fischetti & Jo (2018); Serra et al. (2018). Using these methods, we embed an extra set of variables and constraints in the MILP. This is possible because neural networks are specified as a series of linear inequalities, commonly connected with a rectified linear unit (ReLU) as the activation function. Such a neural network can thus be modeled as a binary mixed-integer linear program, with a continuous variable representing the output of each hidden unit and a binary variable representing the output of each ReLU. Specifically, a single hidden unit in a neural network is represented as:
@@ 4.2 Smart action generation and other computational enhancements
Second, we draw sampled actions according to the implied myopic policy to seed the training process with a set of “reasonable” actions. Specifically, we use the MILP embedding of the Q-network to find the action maximizing the learned approximation of the single-step reward. This allows the method to work out-of-the-box for different settings, without having to explicitly derive a MILP formulation for the optimal myopic action and incur extra expensive computation from repeatedly solving the MIP. We then train the network using temporal difference updates on the sampled actions and rewards to learn their long-term values.
Third, we introduce diversity into the sampled actions with additional random perturbations. One source of diversity is to randomly flip entries of the myopic action (we call this “perturbed myopic”). Another is to randomly sample infeasible actions by, e.g., starting with the all-ones or all-zeros vector for $\bm{a}$ and then randomly flipping a small number of entries (isolating the impact of including or knocking-out individual actions). This leverages the property that, in restless bandits, we can simulate valid state transitions even for infeasible actions because the state transitions are defined independently per-arm. Incorporating perturbed and even infeasible actions greatly increases the diversity of potential samples because for some combinatorial constraints it may be difficult to even find a single feasible solution (and otherwise may be difficult to explore parts of the action space). Similar issues arise if the per-step feasible action space is relatively small (e.g., if the budget is small with $N\gg B$). Including examples of infeasible actions provides a more diverse training set for the Q-network, encouraging better generalization even to feasible actions.
@@ 5 Experiments
To understand the importance of sequential planning in these settings, we consider three myopic baselines. These myopic solution approaches focus on maximizing the expected reward from the immediate next state, based on the current state and transition probabilities (with no Q-network or any type of rollout). We implement a Myopic policy as a MILP that directly encodes (as a linear objective) the expected reward of the immediate next state, while using the same MILP solver as SEQUOIA. Myopic would be optimal in a static (non-sequential) setting, but is not necessarily so in the sequential contexts we consider here. We also use a Sampling approach, which mirrors the method proposed by He et al. (2016), that randomly samples $k=100$ actions of the $\binom{N}{B}$ possible combinatorial actions and picks the action with the largest expected myopic reward. The final myopic baseline is an Iterative Myopic solution construction algorithm that greedily selects one feasible component to build up the full action, similar to Iterative DQN, but using the best single-step reward instead of the predicted Q-value.
@@ 7 Conclusion
We design our experimental settings around restless bandits because they are an increasingly common planning formulation that requires optimizing for long-term reward in a sequential setting over a discrete, combinatorial action space. However, our method is a general approach for stochastic planning with per-timestep combinatorial actions, provided that the feasible set of actions can be described as an MDP. In particular, our use of model-free RL (via DQN) enables direct adaptation to a wide range of planning problems.
@@ E.1 Data generation
Across all experimental settings, the base MDP used to represent each arm is the 4-state MDP described in Section B.2. For each problem setting, we randomly generate transition probabilities and constraints on the actions.

### v2

@@ Abstract
Reinforcement learning (RL) has increasingly been applied to solve real-world planning problems, with progress in handling large state spaces and time horizons. However, a key bottleneck in many domains is that RL methods cannot accommodate large, combinatorially structured action spaces. In such settings, even representing the set of feasible actions at a single step may require a complex discrete optimization formulation. We leverage recent advances in embedding trained neural networks into optimization problems to propose SEQUOIA, an RL algorithm that directly optimizes for long-term reward over the feasible action space. Our approach embeds a Q-network into a mixed-integer program to select a combinatorial action in each timestep. Here, we focus on planning over restless bandits, a class of planning problems which capture many real-world examples of sequential decision making. We introduce coRMAB, a broader class of restless bandits with combinatorial actions that cannot be decoupled across the arms of the restless bandit, requiring direct solving over the joint, exponentially large action space. We empirically validate SEQUOIA on four novel restless bandit problems with combinatorial constraints: multiple interventions, path constraints, bipartite matching, and capacity constraints. Our approach significantly outperforms existing methods—which cannot address sequential planning and combinatorial selection simultaneously—by an average of 24.8% on these difficult instances.
@@ 1 Introduction
Reinforcement learning (RL) has made tremendous progress in recent years to solve a wide range of practical problems Treloar et al. (2020); Marot et al. (2021); Silvestro et al. (2022); Degrave et al. (2022). While successful at dealing with large or infinite state spaces, RL struggles with discrete, combinatorial action spaces. This limitation is pertinent to many real-world sequential decision-making problems, where resource constraints frequently lead to combinatorial action spaces Dulac-Arnold et al. (2020). Consider, for example, a sequential resource allocation problem in which public health workers are dispatched to visit patients. The workers each have a limited daily budget to maximize patient well-being. These requirements give rise to an exponentially large combinatorial action space to optimize over, even when the number of workers and patients is small.
@@
A suitable framework for modeling such problems is that of restless multi-armed bandits Niño-Mora (2023), which have been applied to settings from clinical trial design Villar et al. (2015) to behavioral interventions Mate et al. (2022). An “arm” of the restless bandit corresponds to a patient in the aforementioned public health example. At every timestep, each arm is in one of a finite number of states (e.g., representing the health of that individual), and we aim to select actions so that arms move into the more beneficial states. Typically, restless bandit solutions have been confined to relatively simple and thus tractable problem structures. For example, the common Whittle index policy Whittle (1988) requires that each arm be independent, each action impacts only a single arm, and arms transition independently (Figure 1A). Under these assumptions, the problem admits a threshold-based top-$K$ policy which decouples the combinatorial action selection by learning policies for each arm independently.
@@
Figure 1: Restless bandits are a variant of planning over Markov decision processes, where each “arm” transitions depending on whether it is acted upon. Standard restless bandits can be solved with threshold-based policies, but these approaches are unable to address challenging settings with combinatorial constraints on the actions that cannot be decoupled, prohibiting the application of easy heuristic solutions. Our paper considers this class of strongly coupled restless bandits (coRMAB). We describe these new problem formulations (AD) for restless bandits in detail in Section 3.
@@
Real-world applications, on the other hand, often involve complicated problem structures beyond the simple budget constraint, for which action decomposition would substantially degrade performance. Figure 1A–D illustrate four such scenarios. For example, consider the problem of planning a route for a health worker to visit patients within a travel constraint (see Figure 1D): each action (path) impacts multiple patients and cannot be decomposed further (e.g., into separate edges) while ensuring the overall path is feasible. Another example is where actions represent heterogeneous interventions that may impact not just a single arm but several simultaneously. Individuals may be acted on by multiple actions simultaneously (e.g., exposure to different messaging campaigns or nurse visits), with a potentially nonlinear cumulative effect (Figure 1B).
@@
Here, we consider for the first time sequential combinatorial settings where the reward comes not from a single-shot action but is incurred after enacting a policy over many timesteps. We leverage recent advances in integrating deep learning with mathematical programming, in which a neural network with ReLU activations can be directly embedded into a mixed-integer program Fischetti & Jo (2018); Serra et al. (2018). By integrating this technique with an RL training procedure, we show how to learn the long-term reward using a deep Q-network to approximate the value function, while tractably optimizing this Q-network over the discrete combinatorial space at each step with a mixed-integer program. We call our algorithm SEQUOIA, for SEQUential cOmbInatorial Actions.
@@ 2 Problem description
We consider offline, stochastic planning for restless bandits with combinatorial action constraints, where the transition dynamics and reward are known a priori, but the state and action space are too large to be solved directly. Specifically, we consider the setting where the arms are strongly coupled so we cannot decouple the arms of the bandit into independently controlled Markov chains using a Lagrangian relaxation. We begin by describing the standard restless bandit, then present our general formulation for constrained combinatorial-action RMABs, which we refer to as coRMAB.
@@ 2.2 Restless bandits with strongly coupled actions (coRMAB)
We now introduce coRMABs, a broader class of problems for restless bandits where actions may be strongly coupled due to combinatorial constraints. We consider a set of $N$ actions $a_{i}$ that each impact the transition probability of some subset of the arms. At each timestep, we must pick a combinatorial action over the arms i.e., the action vector $\bm{a}$ must be in $\mathcal{C}\subseteq\mathcal{A}^{\times}$, a set defined by constraints. (We provide example instantiations of $\mathcal{C}$ in Section 3.) Now, the transition probabilities may be coupled, with individual actions impacting multiple arms. We denote the joint transition probability as $P^{\times}:\mathcal{S}^{J}\times\mathcal{A}^{N}\times\mathcal{S}^{J}\rightarrow[0,1]^{J}$.
@@ 4 Solving RMABs with combinatorial actions
We present a novel algorithm, SEQUOIA, which builds on deep Q-learning (DQN) by integrating a mixed-integer linear program (MILP) for combinatorial action selection. We start by introducing the basic DQN+MILP framework in Section 4.1. In Section 4.2, we provide a set of strategies to find a good initialization for the Q-network and improve computational efficiency. We visualize our approach in Figure 2, and Algorithm 1 provides pseudocode for the Q-network training procedure.
While we focus on the restless bandit problem setting, SEQUOIA generalizes to other sequential planning problems with per-timestep combinatorial actions, provided that the restrictions on the actions can be represented as constraints in a mixed-integer program.
@@ 4.1 Q-learning with combinatorial actions
Deep Q-learning aims to estimate the long-run value of action $\bm{a}$ from state $\bm{s}$, denoted $Q(\bm{s},\bm{a})$, by parameterizing the Q-function with a neural network. The DQN algorithm Mnih (2013) samples (action, next state, reward) sequences via a mix of random actions and on-policy samples (i.e., actions that maximize the current estimate of $Q$). The samples are used to improve the Q-function via temporal difference (TD) updates which aim to minimize the residual in the Bellman equation.
@@
Instead, we leverage recent advances that enable encoding a neural network into a mixed-integer linear program to enable optimization over the trained network Fischetti & Jo (2018); Serra et al. (2018). Using these methods, we embed an extra set of variables and constraints in the MILP. This is possible because neural networks are specified as a series of linear inequalities, commonly connected with a rectified linear unit (ReLU) as the activation function. Such a neural network can thus be modeled as a binary mixed-integer linear program, with a continuous variable representing the output of each hidden unit and a binary variable representing the output of each ReLU. Specifically, a single hidden unit in a neural network is represented as:
@@ 4.2 Smart action generation and other computational enhancements
Second, we draw sampled actions according to the implied myopic policy to seed the training process with a set of “reasonable” actions. Specifically, we use the MILP embedding of the Q-network to find the action maximizing the learned approximation of the single-step reward. This allows the method to work out-of-the-box for different settings, without having to explicitly derive a MILP formulation for the optimal myopic action and incur extra expensive computation from repeatedly solving the MILP. We then train the network using temporal difference updates on the sampled actions and rewards to learn their long-term values.
Third, we introduce diversity into the sampled actions with additional random perturbations. One source of diversity is to randomly flip entries of the myopic action (we call this “perturbed myopic”). Another is to randomly sample infeasible actions by, e.g., starting with the all-ones or all-zeros vector for $\bm{a}$ and then randomly flipping a small number of entries (isolating the impact of including or knocking-out individual actions). This leverages the property that, in restless bandits, we can simulate valid state transitions even for infeasible actions because the state transitions are defined independently per-arm. Incorporating perturbed and even infeasible actions greatly increases the diversity of potential samples because for some combinatorial constraints it may be difficult to even find a single feasible solution (and otherwise may be difficult to explore parts of the action space). Similar issues arise if the per-step feasible action space is relatively small (e.g., if the budget is small with $B\ll N$). Including examples of infeasible actions provides a more diverse training set for the Q-network, encouraging better generalization even to feasible actions.
@@ 5 Experiments
To understand the importance of sequential planning in these settings, we consider three myopic baselines. These myopic solution approaches focus on maximizing the expected reward from the immediate next state, based on the current state and transition probabilities (with no Q-network or any type of rollout). We implement a Myopic policy as a MILP that directly encodes (as a linear objective) the expected reward of the immediate next state, while using the same MILP solver as SEQUOIA. Myopic would be optimal in a static (non-sequential) setting, but is not necessarily so in the sequential contexts we consider here. We also use a Sampling approach, which mirrors the method proposed by He et al. (2016), that randomly samples $k=100$ actions of the up to $\binom{N}{B}$ possible combinatorial actions and picks the action with the largest expected myopic reward. The final myopic baseline is an Iterative Myopic solution construction algorithm that greedily selects one feasible component to build up the full action, similar to Iterative DQN, but using the best single-step reward instead of the predicted Q-value.
@@ 7 Conclusion
We design our experimental settings around restless bandits because they are an increasingly common planning formulation that requires optimizing for long-term reward in a sequential setting over a discrete, combinatorial action space. However, our method is a general approach for stochastic planning with per-timestep combinatorial actions, provided that the feasible set of actions can be described as an MDP. In particular, our use of model-free RL (via DQN) enables direct adaptation to a wide range of planning problems. Extending SEQUOIA to broader RL planning problems beyond restless bandits would benefit from more efficient methods to explore the large set of state–action transitions, as well as more structured learning methods, such as graph neural networks or other specialized architectures to capture characteristics of the problem setting.
@@ E.1 Data generation
Across all experimental settings, the base MDP used to represent each arm is the 4-state MDP described in Section B.2. We randomly generate transition probabilities. The constraints on the actions are generated for each problem setting as described in Appendix C.